# From the CDC to emerging infectious disease publics: The long-now of polarizing and complex health crises

### v2

@@ 3. Dataset
We used the twarc Python library for data mining via the Twitter API to collect the tweets used in this study. This was possible through the use of an individualized bearer token granted by Twitter. The inclusion criteria for the tweets in this study were tweets that were produced by users in the United States, in English, posted between January 2020 and January 2022, and not retweets. Furthermore, we specifically only included tweets that were sent to the official @CDCgov account, from @CDCgov, used the hashtag ‘#CDC’, tagged @CDCgov, or mentioned the string ‘CDC’. In total, we collected 21,843 tweets sent to @CDCgov, 11,810 from @CDCgov, 20,012 using ‘#CDC’, 53,104 tagging @CDCgov, and 162,295 mentioning the string ‘CDC’. Finally, we collected any other Tweets using the same “Conversation ID” which indicate that they are under the same discussion thread. In total, 275,124 tweets were collected. While we did not collect all retweets, we do have the number of times each tweet is retweeted with a total of 3,331,350 retweets.
@@ 4. Methods
In this study, we analyze the discourse centered around the CDC using mixed methods. First, we determined the credibility of sources shared on Twitter in §4.1.4. Next, we explored different themes of the discourse using topic models (§4.1.1) and qualitative analysis (§4.1.2). Finally, to better understand the stakeholders and their interests, we clustered users based on their topics of interest and their behavior on the platform (§4.3.1).
@@ 4.1. Features used in data analysis
In this section, we present all the features used to train classifiers, regression, clustering and temporal models we used in our analysis. We start with a description of topics and themes created using topic modeling. We then describe how we measured sentiment, identified low-credibility sources, and finally used platform signals (e.g., retweets) and user properties (e.g., verified user).
@@ 4.1.1. Topic Modeling with BERTopic (N=71)
4.1.2. Qualitative Analysis
In a virtual public square, discourse constructs social contexts (Schiffrin, 1987; Howarth, 2000) as it reflects communication emphasis as expressed by participants (Potter and Wetherell, 1987). When qualitatively analyzing discourse, communication content becomes more than information exchange (i.e., mentions) or diffusion (i.e., retweets) (Matassi and Boczkowski, 2023). Discourse can become a resource to structure knowledge and make sense of a situation (Hardy et al., 2005). Topic models must be interpreted by humans (Chang et al., 2009; Dou et al., 2011) to contextualize them in the social/political environment (Baden et al., 2020; Yarchi et al., 2021). Once the 71 topics with the highest coherence scores were identified, we used qualitative methods to describe each topic and create overarching themes of the issues shared between topic groups as described in Table 15. The following are the steps used in this process:
  1. (1)
    First iteration - describing the top themes: We sampled 100 tweets randomly from each topic to create a general outline of themes that represent frames in our dataset.
  2. (2)
    Selecting top positive, negative, and neutral sentiment comments for each theme: All of the themes had a sentiment distribution with more extreme scores than a normal distribution (Kallner, 2017). Given this distribution, we use log likelihood ratio (LLR) and Tag2Vec to develop top words describing different sentiments associated with each theme. These are presented in Tables 16, 17 and 18. LLR determines whether terms have “topic signatures” that “define the aboutness” (Gupta et al., 2007) of two different sets of words. It was used earlier by Chancelor et al (Chancellor et al., 2018) to determine normative differences between online communities. Similarly, Tag2Vec, using embeddings (similar to the ones discussed in §4.1.1), provides a list of the words closest to a tag (in this case, belonging to a cluster or being tagged as a positive/negative framing of a theme) (Chen et al., 2017). It was used before to determine the linguistic differences between throwaway and pseudonymous users on Reddit (Ammari et al., 2019). We randomly sampled 100 tweets for each category of top theme X sentiment. When there was no sample across any of the sentiment criteria, we sampled another 100 tweets from this theme. This gave us a total of 6,000 tweets.
  3. (3)
    Consensus on themes: After an iterative coding process and weekly meetings to discuss individual code books, we reached consensus on the different themes and the polarization within each (McDonald et al., 2019). We also identified overarching themes as shown in Figure 2.
A note on quoting tweets and identifying Twitter accounts
Fiesler and Proferes (Fiesler and Proferes, 2018) argue that Twitter users do not expect to be quoted verbatim in academic research, even if they are not named. They suggest engaging in Bruckman’s (Bruckman, 2002) recommended levels of user disguise when quoting users in a research study. To protect the privacy of Twitter users, we primarily describe the contents of tweets without direct quotations. We only name public organizations such as the CDC or news organizations and organizations.
Figure 2. Topic groupings into themes
@@ 4.1.3. Sentiment analysis (N=1)
To understand how the sentiments of tweets affected the way that people discussed topics in the CDC public, we used VADER, a sentiment analysis tool that is frequently applied to Twitter text (Elbagir and Yang, 2019) which calculates the valence and polarity for each word (positive, negative, or neutral) along with the percentage of said polarity in a composite score. VADER labels text as having positive sentiment if the composite score is above 0.05, negative sentiment if the score is below -0.05, and neutral sentiment if the score is between -0.05 and 0.05 (Hutto and Gilbert, 2014). We used composite score in regression models, and we used positive, negative, and neutral labels for classifiers (see §4.2).
@@ 4.1.5. Twitter Platform Signals (N=9)
As discussed in §2.1, Twitter provides affordances for propagating messages and expanding your audience. We will use the features identified in the section and listed below:
@@
Additionally, we add the following features to describe the use of Twitter in the CDC public:
  1. (1)
    Total number of Tweets by user (count): this feature represents the total number of Tweets for a user over the period of data collection within the CDC public. The higher this number, the higher the commitment of the user to this discourse.
  2. (2)
    Tenure in the CDC Twitter publics (date of last Tweet - date of first Tweet). While the Tweet count provides one measure of commitment to the discourse, this is another measure that shows how long the user was engaged in this discussion.
  3. (3)
    FFR: Following/Follower ratio: the higher this ratio, the higher the number of people interested in their content without the target user following many others. If the value is less than one, then the user can be described as a “mass follower” who follows other accounts for the“sole purpose of gaining more users” for themselves (Anger and Kittl, 2011).
4.2. Understanding platform signals and content analysis
To better understand platform signals, we trained four regressions to measure what increases the likelihood of retweet counts, quote tweet counts, replies to tweets, and likes for Tweets (see Tables 2 and 3).
Additionally, we trained seven classifiers to better understand what exemplified verified accounts, Tweets responding to the CDC, Tweets using low-credibility links, Tweets using rich media (e.g., pictures and videos), positive, negative, and neutral Tweets based on Vader sentiment analysis (see §4.1.3). The top classifiers, along with the positive and negatively predictive features are shared in Tables 5 and 6.
4.2.1. Hyperparameter Tuning
When building all of the models identified above, we used 82 features as described below:
  1. (1)
    Twitter topics as defined in §4.1.1 (N=71).
  2. (2)
    Vader sentiment analysis as defined in §4.1.3 (N=1).
  3. (3)
    Low credibility source identifier as defined in §4.1.4 (N=1).
  4. (4)
    Twitter platform signals as defined in §4.1.5 (N=9).
We set an 80/20 random train/test split and used Sklearn’s standard scaler to scale data before training (Pedregosa et al., 2011a). To find the top R2 in our regression models, we used Random Search Cross Validation (with 5-fold cross-validation) (Vishnu et al., 2023) to investigate the performance of Ridge and Lasso regressions. We tuned for $\alpha$ and the solver ([”auto”, ”svd”, ”cholesky”, ”lsqr”, ”sparse_cg”, ”sag”, ”saga”, ”lbfgs”]) in Ridge. We also tuned for $\alpha$ and selection ([”cyclic”, ”random”]) in Lasso regression (Wang et al., 2011).
As for the seven classifiers, we again set an 80/20 random train/test split and used Sklearn’s standard scaler to scale data before training. To find the model with the top Area Under the Curve (AUC) - a common metric that evaluates a model’s fit (Swets, 2014; Fawcett, 2006; Lemeshow and Hosmer Jr, 1982) - we used Grid Search Cross Validation (with 5-fold cross-validation) to find the model with the best AUC. A model is considered acceptable as long as the AUC value is above 0.7. The accuracy, recall, F1, and AUC metrics are reported in average and for the top model in Tables 4, 5, and 6. We trained our models on the following classifiers: (1) logistic regression classifier; (2) Random Forest Classifier; (3) Gradient Boosting Classifier; and (4) Gaussian Classifier. They full list of hyperparameters used in the training process can be found in the Appendix (§Classifier hyperparameters).
4.2.2. Explainabilty
For the Ridge and Lasso Regressions, we used the Regressors package (Haas, [n. d.]) to determine significant coefficients (presented in Tables 2 and 3).
However, given that six of our top classifiers were random forest classifiers, whilst the seventh is a Gradient Boosting Classifier, we applied post-hoc analysis to explain the models. We used SHapely Additive exPlanations (SHAP), a post-hoc explainable AI analysis tool to explain each model (Huang and Huang, 2023; Shoemaker et al., 2023). Shapley values, based on game theory (Lundberg and Lee, 2017), allow us to analyze the effects of each feature on output of the model. In turn, this allowed us to rank top predictors and to show dependence between predictors.
4.2.3. Analyzing mention ties: Exponential Random Graph Models
Exponential Random Graph Models (ERGM) model patterns within a network which can be seen as evidence for ongoing structural processes, which are also influenced by node attributes (Robins and Lusher, 2012). ERGM models “extend the logic of multivariate regression to relational data: the presence or absence of a binary tie in a network is an outcome variable predicted by a vector of independent variables called parameters.” (Keegan et al., 2012)[p.431] A positive parameter increases the likelihood of a tie in the network, while a negative parameter indicates a decreased likelihood of tie formation. The parameter equivalent to the intercept in a regression model is the edges parameter which “reflects the log-likelihood of a network tie appearing entirely by chance” which in turn describes the “network density if no other effects were present.” (Keegan et al., 2012)[p.432]. Another network-specific or “endogenous” ERGM parameter is mutuality, or reciprocity which measures the tendency for a tie from one node to another to be reciprocated in the other direction. A higher mutual value increases the value of “cooperative network ties“ between network nodes (Axelrod and Hamilton, 1981). Exogenous node-level coefficients indicate whether actors with a specific attribute are more or less likely to be connected to alters in the network (Lusher et al., 2013).
The method has been used in HCI and CSCW to describe the dynamics of social networks. Keegan et al. (Keegan et al., 2012) built a bipartite network connecting Wikipedia articles and editors who worked on them, finding that the likelihood of an editor contributing to the Wikipedia article increases with every fatality associated with the incident, and where incidents occurred in OECD countries (Keegan et al., 2012)[p.432]. In the medical field, Srinivasan and Jiang (Srinivasan and Jiang, 2023) created a disease co-occurrence network and showed that chronic diseases increase the likelihood of a tie to COVID diagnosis, thus proving a link between the two.
ERGM parameters
First, we will describe the network we formed to analyze using ERGM. Twitter users can explicitly “mention” other users by including their usernames in the text of the tweet. A user u, can mention another user v, by tweeting “@v, what are you up to today.” We created a directed mentions network based on user mentions composed of 59,304 nodes and 315,495 edges. We then used the Statnet R package (Handcock et al., 2019) to fit the ERGM model on the mentions network using the following parameters: (1) Endogenous parameters: (a) edges; (b) mutuality and (2) exogenous parameters: (a) average theme weights as defined in §4.1.2 given the computational cost of incorporating all 71 topics (Kim, 2025); (b) user sentiments broken by positive, negative, and neutral based on the VADER score Positive, negative, and neutral (see §4.1.3); (c) credibility of sources as presented in 4.1.4; and (d) user-specific features as presented in §4.1.5.
@@ 4.3.2. Analyzing cluster content
For each cluster, LLR and Tag2Vec were used to determine the signature of the biographies of cluster users. We also used the top topics discussed by each cluster to guide our qualitative analysis. For larger clusters, 100 users were selected (including at least 10% verified users). For clusters with less than 100 users, all users were qualitatively analyzed. The tweets of each user were read along with their biographies after dividing the samples between team members.
As we did for the topics and themes, the team reached consensus on the description of different clusters after regular team meetings following individual coding of each cluster and sharing memos about their findings with the team (McDonald et al., 2019).
@@ 4.4. Temporal Analysis
Earlier work also analyzed the temporal dynamics of online discourses and how they are affected by the onset of COVID (Lee et al., 2021) by specifying one breaking point, March 11th, which is when COVID was declared a pandemic. The authors then used the Chow test (Wooldridge, 2019) to measure the significance of the change in topic-level discussions before and after that date. However, given that it is difficult to define the different phases of a health crisis beforehand (see §2.3.3), we describe the change-point detection method we used in §4.4.1.
While change-point detection allows us to describe temporal change at the population level, we also wanted to better understand the temporal patterns of use on a personal level. Similar to the study by Gao et al. (Gao et al., 2021), we analyzed users’ commitment to engagement in the online community over time. We describe our survival analysis in §4.4.2.
@@ 4.4.1. Temporal Change-point Detection
We applied the Pruned Exact Linear Time (PELT) change-point detection algorithm (Truong et al., 2020) to identify significant changes in Tweet sentiment and the weight of different topics and their parent themes (as described earlier in §4.1.1).
Change-point detection algorithms find points in a time series where there is a significant change (Killick et al., 2012). We chose the PELT algorithm for change-point detection because it is more conservative in detecting temporal changes, which produces more accurate results (Wambui et al., 2015; Liu et al., 2013). Earlier work adopted this method to analyze Twitter discourse about COVID over time (Valdez et al., 2020).
Once we identified a change-point, we wanted to determine the trend direction (upward or downward trajectory) and significance. To do so, we relied on the Mann-Kendall non-parametric test as presented by Hussain & Mahmud (Hussain and Mahmud, 2019). In figures showing the change-points (e.g., Figures 6, 8, and 10), we used a green spot to identify significant trends, and a blue arrow showing the direction of the trend.
To describe the difference before and after a significant change-point, we used log-likelihood ratio (LLR) to determine “topic signatures” before and after the change-point (Dunning, 1993; Lin and Hovy, 2000) based on the “aboutness” of word lists (Gupta et al., 2007) before and after the change-point (Lee et al., 2021).
4.4.2. Survival Analysis
Following the method used in Gao et al.(Gao et al., 2021), we define user commitment as the duration of engagement in discourse with the CDC in units of participant-month. In our case, the control variables are: (1) retweet count; (2) reply count; (3) like count; (4) quote count; (5) ratio of low-quality links used; (6) posting pictures; and the (7) user being verified. Our independent variables measure engagement in the five main themes identified in Figure 2, as well as average VADER sentiment score for the user.
All continuous variables (other than verified) are standardized using the Sklearn standard scaler (Pedregosa et al., 2011b). We used the lifelines Python package for our analysis (Davidson-Pilon, 2019).
Following the method adopted by Gao et al. (Gao et al., 2021), we used the timestamp for the first Tweet as the starting date for participating in the public sphere. The date of the last Tweet was used as the date of the last activity. Participation is considered censored if the last Tweet is dated from the month the data collection was performed. This is because we do not know when they stop contributing in the public sphere. If a user does not Tweet within three months, we consider them as having dropped out. The model we trained had a concordance of 0.612, which is an acceptable prediction for survival analysis (Zhang et al., 2022b). A summary of the model output is presented in Table 10.
Since the continuous variables are standardized, a hazard ratio (HR) larger than 1 means there is a higher than average likelihood of not surviving (dropping out of the CDC discussions). A hazard ratio less than 1 shows a likelihood of continuing to engage in the CDC public sphere.
5. Study I: Platform signals and message properties
We start this section by analyzing how CDC discourse is affected by Twitter platform affordances (e.g., retweets, likes, etc.) in §5.1.1. Next, we analyze how credibility of sources in a Twitter message, the use of rich media, and the sentiment of that message affect COVID discourse in the CDC public in §5.2. Finally, in §5.3, we analyze how user properties, like being verified or influential in the network (based on FFR), are associated with user behavior in the CDC public.
Summary of findings:
Discussions about vaccines are more likely to receive all kinds of Twitter engagement. Although positive tweets discussed the seriousness of COVID-19, they also shared news about dropping infection levels, especially after the lockdowns. Neutral accounts were more likely to discuss the lack of clarity in the CDC guidelines but avoided talking about mortality rates or engagement in political discourse–both features predictive of negative Tweets. We found that verified accounts, while less likely to Tweet, are more likely to be retweeted (thus the high echo score in Table 1), liked, receive responses, but not quote-tweets, and to use rich media. Tweets using low-credibility sources are also more likely to use rich media and be posted by verified accounts. The more frequently an account uses low-credibility sources, the more likely they are to mention other users.
Table 1. This table compares the CDC, responses to it, verified users, and responses to them, tweets using pictures and responses to them, as well as the use of low-credibility sources and responses to these tweets. Specifically, we focus on the echo and average sentiment for each. Echo is the ratio between retweet/tweets. The closer the value to 1, the less the likelihood of their tweets being retweeted. The higher the echo the higher the number of retweets per Tweet. At times, the properties might overlap. For example, we show how often verified users use images.
Behavior and Affordances Total Unique Echo Sent. Low-cred. (%) Rich (%) Ver. (%)
CDC 2,688,505 11,810 227.65 0.17 0 100 100
Respond to CDC 122,813 21,843 5.623 -0.01 4.16 34.50 6.12
Verified 2,903,733 26,428 109.87 0.10 0.10 75.61 -
Respond to verified 139,058 26,205 5.31 -0.01 3.96 35.07 11.02
Rich media 3,032,625 99,004 30.63 0.01 11.54 - 20.18
Respond to pictures 136,131 27,783 4.90 -0.01 4.28 43.98 8.07
Low credibility 36,008 17,433 2.07 -0.16 - 65.55 5.11
Respond to low credibility 702 468 1.50 -0.07 62.61 61.11 2.78
5.1. Platform affordances, user properties, sentiment and information credibility
We start this section by analyzing how Twitter engagements afforded by the platform like retweets, quote-tweets, likes, mentions, and responses are associated with different themes of COVID discussions. We will specifically focus on direct responses to the CDC and, in turn, direct responses to users from the CDC (as opposed to broadcast messaging).
5.1.1. What is in a retweet and what gets more likes?
As Table 2 shows, discussions for or against the vaccine are more likely to be retweeted and to receive likes. While discussions of lock-downs and re-openings, especially in schools (Topics 38 and 65), were less likely to receive likes, they were retweeted at a higher rate. Similarly, engaging in Topic 1, introducing data showing that COVID is not a hoax, increases the chance of being retweeted while reducing the chance of receiving a like.
The chance of receiving likes and being retweeted is higher when Tweets are posted by verified users and when there is a high number of Tweets in the CDC public. Users with a higher follower-following ratio (FFR) and more direct replies are also more likely to receive likes. However, the more direct replies received by a user, the less likely they are to be retweeted. We discuss direct replies next.
5.1.2. What gets more direct replies and quote-Tweets?
Engaging in discourse for and against the vaccine (Topic 0) increases the chances of receiving a response and being quote-tweeted. The same applies to discussing masking in schools (Topic 26). Both of these topics were polarizing as they pit people who believed that each mitigation tool was important to control COVID spread and others who were against their implementation because (1) they thought that the mitigation strategies are not as effective as advertised; or (2) they thought that forcing these mitigation measures were against their freedom. A related topic early in the pandemic focused on the very presence of COVID as a virus. While some were questioning the magnitude of the effects on society, others were sharing data, especially from the Johns Hopkins dataset, to counter data from the CDC and other official channels that were considered not forthcoming about the effects of the virus. However, Topic 38, focusing on the effects of lock-downs on society, received fewer responses and quote-tweets. Interestingly, this topic received fewer likes as well. However, it did increase the number of retweets.
While being verified is more associated with being retweeted, being a verified user makes it less likely to be quote-tweeted. Furthermore, quote-tweeted users have a lower follower-following ratio (influence in the network) and less Tweets in the CDC public. Next, we focus on analyzing direct responses to the CDC.
Responses to the CDC
As presented in Table 4, the use of rich content was more predictive of responses to the CDC. Much of the rich content included what users referred to as “receipts,” screenshots of earlier CDC messages showing the evolving masking guidelines after vaccinations (Topic 4). Other top topics in responses include the vaccines (Topic 0), the role of the CDC in evictions post-COVID, discussing how COVID is airborne (Topic 36), and comparing the response to COVID with the response to the H1N1 breakout under Obama’s administration (Topic 56).
The framing of these responses went in one of three directions: (1) arguing that the CDC’s guidelines are not strict enough; (2) arguing that the CDC is introducing mandates without the right to do so; (3) the CDC is overreacting since the effects of COVID are overblown (see Figure 20).
Replies to the CDC were less likely to receive direct replies, likes, retweets, quote-tweets, and usually had lower follower-following ratios.
Direct responses from the CDC
We found only 58 direct responses from the official CDC account to 25 accounts based on a question or suggestion. The responses revolved around providing information about the shingles or flu vaccines. A few others were answering questions about the danger of contracting COVID-19 by touching products imported from China. The @CDCgov account answered those accounts by saying there is no danger in handling products from China.
Now that we analyzed responses to and from the CDC, we will next analyze mention co-occurrence networks along with the @CDCgov account.
Table 2. This Table shows top regression results and the top predictors for Retweets ($\alpha=0.31$, solver=svd) and Likes ($\alpha=0.34$, solver=svd). Table shows significant predictors from regression model. **** p¡0.0001; *** p¡0.001; ** p¡0.01; * p¡0.05
Retweets Ridge $R^{2}$=0.90 Likes Ridge $R^{2}$=0.83
Feature Coefficient p-value Feature Coefficient p-value
Topic 0 0.611 **** Topic 0 3.511 ****
Topic 1 0.101 * Topic 1 -0.614 *
Topic 38 0.072 ** Topic 38 -0.432 **
Topic 65 0.045 * Topic 65 -0.23 *
Verified 2.329 * Verified 15.033 ****
like_count 2838.533 **** quote_count -33146.7 ****
quote_count 7534.939 **** reply_count 1189.377 ****
reply_count -1084.07 **** retweet_count 47676.71 ****
count 0.412 **** count 8.882 ****
FFR 17.385 *
Table 3. This Table shows top regression results and the top predictors for responses ($\alpha$=0.31, solver=svd) and quote-tweets ($\alpha$=0.68, solver=cholesky), Quotes-tweets, and Responses. Table shows significant predictors from regression model. **** p¡0.0001; *** p¡0.001; ** p¡0.01; * p¡0.05
Responses Ridge $R^{2}$=0.64 Quote-tweets Ridge $R^{2}$=0.70
Feature Coefficient p-value Feature Coefficient p-value
Topic 0 0.161 **** Topic 0 0.008 ****
Topic 1 -0.06 * Topic 15 -0.008 **
Topic 15 0.036 **** Topic 26 0.011 *
Topic 26 0.004 **** Topic 38 -0.005 *
Topic 38 -0.001 **** Verified -0.032 ****
Verified 0.732 **** reply_count 856.81 ****
like_count 32.959 **** retweet_count 962.706 ****
quote_count 2964.705 **** count -1.813 ****
retweet_count -504.582 **** FFR -0.666 ****
count 10.537 ****
FFR 2.828 ****
Table 4. Predictors for direct replies to the CDC. The table shows the average metrics from RandomSearch and the hyperparameters for the top model. They also show the top positive and negative predictors as determined using SHAP.
Reply to CDC
Average Metrics
Accuracy 0.645
Precision 0.626
Recall 0.767
F1 0.684
AUC 0.716
Top Model (GBC) est=150, max. depth=5, learning rate=0.3
Accuracy 0.752
Precision 0.725
Recall 0.811
F1 0.766
AUC 0.837
Positive Features 0, 4, 27, 36,56, Rich
Negative Features Rep, Like, FFR, RT, 7, 15, 15, 61, QT, 29, 9, 19, 8, 56
Table 5. Part 1 of Classifiers: Performance metrics for Positive, Neutral, Negative, and Reply to CDC classifiers. The table shows the average metrics from RandomSearch and the hyperparameters for the top model. They also show the top positive and negative predictors as determined using SHAP.
Positive Neutral Negative
Average Metrics
Accuracy 0.668 0.681 0.688
Precision 0.667 0.684 0.715
Recall 0.717 0.713 0.593
F1 0.685 0.689 0.633
AUC 0.700 0.719 0.727
Top Model (RF) est=50; mins=5, max=None est=100; mins=5, max=10 est=100; mins=5, max=10
Accuracy 0.984 0.983 0.985
Precision 0.985 0.983 0.984
Recall 0.982 0.983 0.985
F1 0.984 0.983 0.985
AUC 0.999 0.998 0.999
Positive Features RT, 15, QT, FFR, VR, 1 5, FFR, 21, 14, 9, 32, 63 19, 25
Negative Features 25, Rep, 16, 2, 54, 23, 8, 14, 44,29,56 70, Rep, QT, Like FFR, 14, 30, RT, 21
Table 6. Part 2 of Classifiers: Performance metrics for Photo, Low Quality, and Verified classifiers. The table shows the average metrics from RandomSearch and the hyperparameters for the top model. They also show the top positive and negative predictors as determined using SHAP.
Rich Media Low Credibility Verified
Average Metrics
Accuracy 0.719 0.699 0.830
Precision 0.820 0.708 0.838
Recall 0.529 0.754 0.812
F1 0.619 0.718 0.822
AUC 0.756 0.789 0.905
Top Model (RF) est=100; mins=5, max=10 est=100; min=2, max=None est=100; mins=5, max=10
Accuracy 0.996 0.782 0.894
Precision 0.998 0.803 0.876
Recall 0.994 0.748 0.916
F1 0.996 0.775 0.896
AUC 0.995 0.872 0.964
Positive Features RT, FFR, QT, Like, Rep, 5 19, Rich, 7, VR FFR, QT, RT, Like, Rep, 14, Rich
Negative Features VR, 6, 16, 61 Like, Rep, RT, 23, 32 Count, 21
5.1.3. What is in a mention?
Mentioning accounts can be directing a message to an account expecting a response or increasing attention to the message by mentioning an influential account in the network, in this case, the CDC. The co-occurrence mentions network (Nasim and Rajput, 2017) shown in Figure 4 where the top accounts mentioned alongside the @CDCGov account are presented in the following groups:

### v3

@@ 3. Dataset
We used the Python library Twarc to mine tweets via the Twitter API, authenticated through a personal bearer token granted by Twitter. Our dataset includes tweets that met the following criteria: authored in English, originating from U.S.-based users, posted between January 2020 and January 2022, and not classified as retweets. We focused on tweets directly engaging with the Centers for Disease Control and Prevention (CDC) by including messages that were: (1) sent to @CDCgov, (2) posted by @CDCgov, (3) used the hashtag #CDC, (4) tagged @CDCgov, or (5) contained the string “CDC” in the tweet text. This resulted in the collection of 21,843 tweets sent to @CDCgov, 11,810 from @CDCgov, 20,012 with the hashtag #CDC, 53,104 tagging @CDCgov, and 162,295 mentioning “CDC.”
Additionally, we included any tweets sharing the same Conversation ID, indicating participation in the same discussion thread. This yielded a final dataset of 275,124 unique tweets. Although full retweet content was not collected, we retained retweet counts, totaling 3,331,350 retweets across the dataset.
@@ 4. Methods
In this section, we describe the mixed-methods approach used to analyze CDC-related discourse on Twitter. We detail our dataset construction, feature extraction, and topic modeling techniques, followed by sentiment and credibility analysis. We then outline our clustering method to identify discourse communities, our interpretive strategies for qualitative analysis, and the temporal techniques used to detect shifts in public attention over time.
@@ 4.1. Features used in data analysis
In this section, we describe the features used to cluster distinct publics and to examine how discourse evolved both within these clusters and across the broader conversation over time. We begin with an overview of the topics and themes generated through topic modeling, followed by our approach to measuring sentiment and identifying low-credibility sources. We then outline additional platform signals (e.g., retweets) and user attributes (e.g., verification status) incorporated into the analysis.
@@ 4.1.1. Topic Modeling with BERTopic (N=71)
Figure 2. BERTopic clustering results grouped into five major themes, including COVID mitigation, partisan discourse, and public health equity. Themes reflect how discussions clustered semantically and ideologically.
@@ 4.1.2. Sentiment analysis (N=1)
To understand how the sentiments of tweets affected the way that people discussed topics in the CDC public, we used VADER, a sentiment analysis tool that is frequently applied to Twitter text (Elbagir and Yang, 2019) which calculates the valence and polarity for each word (positive, negative, or neutral) along with the percentage of said polarity in a composite score. VADER labels text as having positive sentiment if the composite score is above 0.05, negative sentiment if the score is below -0.05, and neutral sentiment if the score is between -0.05 and 0.05 (Hutto and Gilbert, 2014).
@@ 4.1.4. Twitter Platform Signals (N=9)
Twitter provides affordances for propagating messages and expanding your audience. We will use the features identified in the section and listed below:
@@ 4.2.2. Analyzing cluster content
To interpret discourse within each cluster, we combined computational and qualitative approaches rooted in discourse analysis. We used two methods to characterize the linguistic patterns of cluster users: Log Likelihood Ratio (LLR), which identifies distinguishing terms to define the “aboutness” of a discourse (Gupta et al., 2007), and Tag2Vec, which uses document embeddings to find semantically similar words associated with cluster tags or sentiment framings (Chen et al., 2017). These methods helped us surface salient terms in both tweet content and user biographies.
Our qualitative analysis followed a multi-step protocol. For large clusters, we sampled 100 users, including at least 10
Drawing from discourse theorists who argue that language does more than transmit information—it also shapes social realities (vanDijk1985semantic; Schiffrin, 1987; Potter and Wetherell, 1987)—we treated tweets not merely as retweet or mention events but as meaningful communicative acts (Papacharissi, 2016; Utz et al., 2013). In this view, discourse becomes a resource through which users structure knowledge and make sense of uncertainty (vanDijk1985semantic). Because topic models are not self-interpreting, we applied qualitative interpretive strategies to contextualize each of the 71 topics generated with the highest coherence scores (Chang et al., 2009; Dou et al., 2011; Park and Lee, 2020; Chancellor et al., 2018).
This interpretive process involved three major steps:
  1. (1)
    Theme Identification: We randomly sampled 100 tweets from each topic to draft preliminary framings that reflected key thematic patterns.
  2. (2)
    Sentiment Differentiation: Using LLR and Tag2Vec, we identified the top words associated with each theme by sentiment polarity (positive, negative, neutral). We sampled 100 tweets for each theme–sentiment pairing, generating a corpus of 6,000 tweets. Sentiment distributions were wide-tailed and highly polarized, consistent with prior work showing non-normal emotional distributions in crisis discourse (Kallner, 2017).
  3. (3)
    Consensus Coding: Team members collaboratively coded tweets and developed shared analytic memos. We discussed interpretations in weekly meetings to refine categories and identify overarching themes and tensions (McDonald et al., 2019).
In line with best practices in social media research ethics, we do not quote individual users verbatim unless they represent public institutions such as the CDC or major news outlets. This follows guidance from Fiesler and Proferes (Fiesler and Proferes, 2018), who argue that users do not expect academic reuse of their tweets, even if technically public. Instead, we use paraphrased examples and thematic summaries, consistent with Bruckman’s framework for protecting user anonymity online (Bruckman, 2002).
4.2.3. Analyzing cluster structure
We used Krackhardt and Stern’s (Krackhardt and Stern, 1988) “E-I” index —a measure of network insularity of cohesion. The index ranges between [-1,1]. The closer the value to +1, the more externally connected the network while the closer to -1, the higher the chance of an echo chamber being formed in the network (Wolfowicz et al., 2023; Bruns, 2017).
The ”E-I” index was built on stance (topic X sentiment) network built using mentions and responses as edges (reflecting the three propagation affordances in Twitter).
@@ 4.3. Temporal Analysis
Earlier work also analyzed the temporal dynamics of online discourses and how they are affected by the onset of COVID (Lee et al., 2021) by specifying one breaking point, March 11th, which is when COVID was declared a pandemic. The authors then used the Chow test (Wooldridge, 2019) to measure the significance of the change in topic-level discussions before and after that date. However, given that it is difficult to define the different phases of a health crisis beforehand (see §LABEL:related_work:flashpoint), we describe the change-point detection method we used in §4.3.1.
@@ 4.3.1. Temporal Change-point Detection
We used the Pruned Exact Linear Time (PELT) algorithm for change-point detection to identify significant shifts in sentiment and topic prevalence over time within each discourse cluster. PELT is known for its conservative and accurate detection of temporal changes in time series data (Truong et al., 2020; Killick et al., 2012; Wambui et al., 2015), and has been applied in prior Twitter analyses, such as COVID-19 discourse (Valdez et al., 2020).
To interpret the direction and significance of trends between change-points, we applied the Mann-Kendall non-parametric trend test (Hussain and Mahmud, 2019). In visualizations, red dashed lines indicate change-points; arrows show trend direction; and green dots mark statistically significant segments (p ¡ 0.05).
5. Findings
In this section, we present our key findings across five thematic areas. We begin by identifying patterns of echo chambers and cross-cutting discourse within the CDC Twitter publics. We then analyze how users responded to CDC mitigation guidelines, including shifts in messaging clarity and polarization. Beyond COVID, we examine discourse around other CDC guidelines, explore the politicization of public health communication, and track how marginalized communities—particularly those focused on racial equity—sustained engagement over time despite declining broader interest. Throughout this section, we analyze the mostly polarized discussions under each theme, and we introduce the different stakeholders and their respective positive, neutral, and negative communication frames.
To get an overview of the discourse, we analyzed the top 10 hashtags and their co-occurrence network with #CDC (Türker and Sulak, 2018) (Figure 4). This revealed three main themes: